Monday 14th July 2014, 14.30 BST.
29 Open UK tickets today. I might have to send my apologies to this week's meeting as Lancaster is receiving a delivery Tuesday morning.
FNAL VOMS TICKETS
As seen on TB-SUPPORT - a number of sites got tickets concerning jobs still contacting the FNAL voms server for CMS/ILC. Birmingham, RHUL, Liverpool and the Tier 1's tickets are still being worked on - RHUL's ticket might not have been spotted yet (still assigned).
DECOMMISSIONING THE FTS3 SERVICE
Gareth opened a ticket to document the retirement, in accordance with ancient grid laws. As naught is happening until the 2nd of September I put on hold till nearer the time. On Hold (14/7)
enmr.eu wanted to add tags to one of the Tier 1's arc ces, which of course didn't work. There was an interesting exchange about why a VO would still want to have a site publish tags in the age of cvmfs (essentially so they can minimise changes to the submission gubbins). Andrew offered to add in the tag "VO-enmr.eu-CVMFS" by hand to his CE, it's likely that other sites might be asked to do the same - and it's a solution worth noting for other VOs. In progress (14/7)
Enabling HyperK at the Tier 1. Ticket looks a little stalled after Chris commented that it was wise for Hyper K to be enabled on only Arc-CEs (in light of RAL going dairy free). In progress (2/7)
UCL are still having trouble with nagios tests after a pool node died. Ben is having trouble getting the new disk server set up - I tried to give him some tips and advised shouting out for help. In progress (8/7)
Bristol having trouble with CMS transfers- Lukasz noticed Storm was being odd (believing there to be no free space when there was). The SE was kicked but the problem (or a similar one) showed up again. Anyone seen similar? (Looking at Chris Walker:Storm Sage again here). In Progress (9/7)
cf TIER 1 ticket https://ggus.eu/index.php?mode=ticket_info&ticket_id=106324
CMS pilots losing contact with their home base. Looks similar to the issue at RAL, where they seem to have had some success (still waiting to see if it was complete). If the RAL chaps could elaborate on the firewall tweaks that brought about this improvement it would be greatly appreciated (The RAL ticket could do with an update too)! In Progress (14/7)